OpenStudy (gabbyalicorn):

What is the exact area of a circle whose diameter is 18 cm?

jimthompson5910 (jim_thompson5910):

formula A = pi*r^2 pi = 3.14 approximately r = radius = half of diameter

OpenStudy (gabbyalicorn):



jimthompson5910 (jim_thompson5910):

ok so each answer is in terms of pi, so ignore the pi's for now

OpenStudy (gabbyalicorn):

whats the radius?

jabez177 (jabez177):

18 divided by 2

jimthompson5910 (jim_thompson5910):

diameter = 18 radius = half of diameter

OpenStudy (gabbyalicorn):

so 9

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

then you plug that into A = pi*r^2

OpenStudy (gabbyalicorn):

3.14 x 18

jabez177 (jabez177):

\[3.14 \times 8 \times 8 \]

OpenStudy (gabbyalicorn):

200.96

OpenStudy (gabbyalicorn):

divide it by 2?

jabez177 (jabez177):

Wait... Let me try this again...

jimthompson5910 (jim_thompson5910):

again, ignore the pi's

jimthompson5910 (jim_thompson5910):

every answer choice has a pi in it

OpenStudy (gabbyalicorn):

8x8 = 64

jimthompson5910 (jim_thompson5910):

but you found the radius to be 9, not 8

OpenStudy (gabbyalicorn):

9 x 8 = 72

jimthompson5910 (jim_thompson5910):

r^2 means r times r

jimthompson5910 (jim_thompson5910):

r = 9

OpenStudy (gabbyalicorn):

9 x8 x8 ?

OpenStudy (gabbyalicorn):

576

jimthompson5910 (jim_thompson5910):

idk how you're getting 9 x 8

jimthompson5910 (jim_thompson5910):

r^2 = r x r = 9 x 9 = ?? I replaced the 'r's with 9 since the radius is 9

OpenStudy (gabbyalicorn):

81?

jimthompson5910 (jim_thompson5910):

yep so 81pi is your answer

OpenStudy (gabbyalicorn):

Thank you!!!!